# RanLayNet: A Dataset for Document Layout Detection used for Domain Adaptation and Generalization

### v1

@@ 2. Related Work
Early methodologies in document layout analysis relied on rule-based algorithms and heuristic approaches (Ahmad et al., 2016). However, the contemporary landscape has witnessed a shift towards addressing this issue through deep learning techniques. They involve the utilization of object detection models (Girshick et al., 2014; Girshick, 2015; Ren et al., 2015; Dai et al., 2021; Jocher et al., 2021; Carion et al., 2020; Tan et al., 2020), which have demonstrated substantial enhancements in accuracy and speed over the past decade. Notably, most cutting-edge object detection methods can be easily trained and implemented with minimal effort due to standardized ground-truth data formats(Lin et al., 2014) and the widespread adoption of common deep learning frameworks (Wu et al., 2019). Reference datasets such as PubLayNet (Zhong et al., 2019) and DocBank (Li et al., 2020) conveniently present their data in the widely accepted COCO format (Lin et al., 2014).

### v2

@@ 1. Introduction
[…]
Doclaynet Results on Fine-tuned RanLayNet model Refer to caption
@@ 2. Related Work
Early methodologies in document layout analysis relied on rule-based algorithms and heuristic approaches (Ahmad et al., 2016). However, the contemporary landscape has witnessed a shift towards addressing this issue through deep learning techniques. They involve the utilization of object detection models (Girshick et al., 2014; Girshick, 2015; Ren et al., 2015; Dai et al., 2021; Jocher et al., 2021; Carion et al., 2020; Tan et al., 2020; Anand et al., 2024), which have demonstrated substantial enhancements in accuracy and speed over the past decade. Notably, most cutting-edge object detection methods can be easily trained and implemented with minimal effort due to standardized ground-truth data formats (Lin et al., 2014) and the widespread adoption of common deep learning frameworks (Wu et al., 2019). Reference datasets such as PubLayNet (Zhong et al., 2019) and DocBank (Li et al., 2020) conveniently present their data in the widely accepted COCO format (Lin et al., 2014).
@@ 3. Datasets
[…]
Pipeline of RanLayNet generation. Crops of PubLayNet are randomly pasted on white canvas on the basis of remaining spaces on the canvas. Refer to caption
@@ 8. Acknowledgement
Appendix A Appendix
Noise labeling, refers to the process of intentionally adding noisy labels to the training data in order to study the effect of noisy labels on the performance of models. This is often done to evaluate the robustness of machine learning algorithms to different types and levels of label noise.
The training/validation curves for YOLOv8 measures on IIIT-AR-13K, PubLayNet, and RanLayNet are in Fig. 4, Fig. 5, and Fig. 6 respectively. The training loss curve depicts improved performance over epochs, signifying effective learning and adaptation.
The validation loss curve evaluates generalization by reducing loss on separate data. Convergence of both curves indicates successful training, reflecting enhanced object detection accuracy. This understanding guided our approach, yielding a proficient model for real-world scenarios.
[…]
YOLOv8 Training/Validation Loss Curves for IIIT-AR-13k Dataset Refer to caption
Figure 4. YOLOv8 Training/Validation Loss Curves for IIIT-AR-13k Dataset
[…]
YOLOv8 Training/Validation Loss Curves for PubLayNet Refer to caption
Figure 5. YOLOv8 Training/Validation Loss Curves for PubLayNet
[…]
YOLOv8 Training/Validation Loss Curves for RanLayNet Refer to caption
Figure 6. YOLOv8 Training/Validation Loss Curves for RanLayNet